OpenStudy (anonymous):

The first five terms of a sequence are shown below: 5, 9, 13, 17, 21 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 5n - 1; n ≥ 1 f(n) = 5n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1

OpenStudy (tkhunny):

"Best"? What a strange word! Since only the last one works AT ALL, we probably should pick it. f(n) = 5n - 1 f(n) = 5n + 1 f(n) = 4n - 1 f(n) = 4n + 1 Substitute n = 1 f(1) = 5(1) - 1 = 5-1 = 4 f(1) = 5(1) + 1 = 5+1 = 6 f(1) = 4(1) - 1 = 4-1 = 3 f(1) = 4(1) + 1 = 4+1 = 5